OpenStudy (anonymous):

solve with substitution method 9x-6y=-24 7x+48=y

OpenStudy (anonymous):

This answer is wrong, where does 2016 come from?

OpenStudy (anonymous):

I get x = -1/8

OpenStudy (anonymous):

9x-42x-288=-24 -33=264 x=-8 y=7*-8+48 y=-8

OpenStudy (anonymous):

no its -8,-8